OpenStudy (anonymous):

Find all solutions to the equation. cos(Theta)-1=-1 Answer needs to be in radians in terms of pi

OpenStudy (anonymous):

is it \[\cos(\theta)-1=-1\]?

jimthompson5910 (jim_thompson5910):

cos(theta) - 1 = -1 cos(theta) - 1+1 = -1+1 cos(theta) = 0 arccos(cos(theta)) = arccos(0) theta = pi/2 + 2pi*n or theta = 3pi/2 + 2pi*n theta = pi/2 + pi*n

OpenStudy (anonymous):

same as \[\cos(\theta)=0\] so \[\theta =\frac{\pi}{2}+k\pi,k\in \mathbb Z\]

OpenStudy (anonymous):

thanks, ya guys:)